to keep the money supply at its original level, does it culate, in dollars, how much the central bank . Explain how banks create money 5. What are the various ways in which the Federal 6. As a Case Study in the chapter discusses, the Reserve can influence the money supply? money supply fell from 1929 to 1933 because Why might a banking crisis lead to a fall in the money supply? of the money supply and the data in Table 4-2 questions TIONS a. What would have happened to the money sup- risen e.The central bank decides to increase the money supply by 10 percent. In each of the above four scenarios, how much should it increase the monetary base? larded . In the nation ofWiknam, people the reserve deposit ratio had remained the same? b. What would have happened to the money sup- ply if the reserve-deposit ratio had risen but the the same? 4· hold $1,000 of currency and $4,000 of demand deposits in the only bank, Wikbank. The reserve-deposit ratio is 0.25 a. What are the money supply, the monetary sible for the fall in the money supply? 7. To increase tax revenue, then Sprernment in 1932 imposed a 2-cent tax on checks written on bank account deposits. (In todays dollars, this tax would amount to about 34 cents per check.) a. How do you think the check tax affected the base, and the money multiplier? b. Assume that Wikbank is a simple bank: it takes in deposits, makes loans, and has no capital Show Wikbanks balance sheet. What value of loans does the bank have outstanding? b. Use the model of the money supply under . Wiknams central bank wants to increase the fractional-reserve banking to discuss how this money supply by 10 percent.Should it buy or tax affected the money supply in open-market opera- tions? Assuming no change in the moncy multiplier, calculate, in dollars, how much central bank needs to transact For any problem marked with LounchPod, there is problem. To access these interactive, step-by-step LounchPod In the economy of Panicia, the monetary base is SI.αο. People hold a third of their moncy in the form of currency (and thus two-thirds as bank deposits). Banks hold a third of their deposits in reserve a. What are the reserve-deposit ratio, the deposit ratio, the money multiplier. and the money supply? b. One day, fear about the banking system strikes the population, and people now want to hold half their money in the form of currency. If the central bank does nothing, what is the new money supply? e. I the face of this panic, the central bank

need an answer to question 5

textbook is macroeconomics 9th edition
Answer #1

(5)

(a)

Reserve-deposit ratio (rr) = 1/3

Currency-deposit ratio (cr) = 1/3

Money multiplier (mm) = (1 + cr) / (cr + rr) = [1 + (1/3)] / [(1/3) + (1/3)] = (4/3) / (2/3) = 2

Money supply (MS) = Monetary base (MB) x mm = $1,000 x 2 = $2,000

(b)

New value of cr = 1/2

mm = [1 + (1/2)] / [(1/2) + (1/3)] = (3/2) / (5/6) = 1.8

New MS = $1,000 x 1.8 = $1,800

(c)

Since money supply has decreased, central bank has to increase money supply using open market purchase of government securities.

Decrease in MS = $2,000 - $1,800 = $200, so central bank has to increase money supply by $200.

Required open market purchase of bonds = Required increase in MS / New mm = $200 / 1.8 = $111.11